OpenStudy (anonymous):

anyone familiar with this: sin(x) - 2sin(x)cos(x) = 0

OpenStudy (anonymous):

You need to factor this and then set each factor equal to 0. So set it up like this: \[\sin(x) - 2\sin(x)\cos(x) = 0 \implies \sin(x)(1-2\cos(x)) = 0\]

OpenStudy (anonymous):

Solve the 2 equations: sinx = 0 1- 2cosx = 0

OpenStudy (anonymous):

hey its you!!

OpenStudy (anonymous):

still having a hard time resolve this i got 0, pi, 2pi and pi/3

OpenStudy (anonymous):

Did your problem give you a restriction? Like did it say \(0 \le\theta \ <\ 2\pi\) or did it not give any restrictions?

OpenStudy (anonymous):

(a) Solve for all radian solutions. Give all answers as exact values in radians. Do not use a calculator. (Enter your answer in the form a + bk, where a is in [0, 2π), b is the smallest possible positive number of radians, and k represents any integer. Enter solutions from smallest to largest. If there are any unused answer boxes, enter NONE in the last boxes.)

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Alrighty then, so all solutions basically. So since it wants [0,2pi), we can kick out your 2pi answer. So you have 0, pi, pi/3. Youre missing one :)

OpenStudy (anonymous):

5pi/3

OpenStudy (anonymous):

??

OpenStudy (anonymous):

yay!! i got it right thank you

OpenStudy (anonymous):

Lol, okay, cool. Good job :)

OpenStudy (anonymous):

well now that your here can you help me with some more??

OpenStudy (anonymous):

Whatcha got?

OpenStudy (anonymous):

2sin^2(theta)-2sin(theta)-1=0 .... i got sin = 1/2, 1

OpenStudy (anonymous):

so 30 degrees, and 90??

OpenStudy (anonymous):

so since sin is positive its btwn 0 and 180

OpenStudy (anonymous):

i put down 30 and 150 and it was wrong...

OpenStudy (anonymous):

Yeah, your answers arent going to be quite that pretty, this equation doesnt factor so nicely. \[2\sin^{2}( \theta ) - 2\sin (\theta) - 1 = 0\] So I'd be plugging these numbers into the quadratic formula to see what we get. \[\frac{ 2 \pm \sqrt{(-2)^{2} - 4(2)(-1)} }{ 2(2) } = \frac{ 2 \pm \sqrt{12} }{ 4 } = \frac{ 2 \pm 2\sqrt{3} }{ 4 } = \frac{ 1 }{ 2 } \pm \frac{ \sqrt{3} }{ 2 }\]

OpenStudy (anonymous):

wow, cant you do (2sin-1) (sin-1)

OpenStudy (anonymous):

That would give you: \[2\sin^{2} ( \theta ) - 3 \sin ( \theta) + 1\] Did you type the correct equation above? Maybe you mistyped it and thats why I have an odd result?

OpenStudy (anonymous):

nope, your right.. ok so I need to use the quadratic equation

OpenStudy (anonymous):

Right. Which of course gives you some awkward answers. \[\sin^{-1} (\frac{ 1 }{ 2 }+\frac{ \sqrt{3} }{ 2 })\] \[\sin^{-1}(\frac{ 1 }{ 2 }- \frac{ \sqrt{3} }{ 2 })\] Now the question is are these both valid answers or no?

OpenStudy (anonymous):

please tell me i dont have to put each of them into an equation

OpenStudy (anonymous):

Well, you wouldnt be able to simplify those answers. I'm just making sure you know why they would or would not be valid answers.

OpenStudy (anonymous):

no i dont think they are valid..

OpenStudy (anonymous):

oh wait they are

OpenStudy (anonymous):

Yes, they are. The domain for sin^(-1)x is [-1,1], so as long as what I plug in is in that interval, they work.

OpenStudy (anonymous):

im suppose to come up with a degree so should i just put it in the calculator?

OpenStudy (anonymous):

I suppose so. Can't really do anything else.

OpenStudy (anonymous):

if it is bigger than 1 then its not possible, right?

OpenStudy (anonymous):

If the input is larger than 1 then yes, its not possible.

OpenStudy (anonymous):

yay! thank you... i might just keep you here like my security blanket... what are you doing awake right now anyways?

OpenStudy (anonymous):

Im often awake at this hour. Unless Im just super tired, its hard for me to be in bed before midnight. Not sure how late ill be up, maybe another hour or two.

OpenStudy (anonymous):

wow! i usually am in bed by now... but Im behind in my hw... and I am trying to avoid normal life... so here I am doing hw

OpenStudy (anonymous):

I always tryto avoid normal life, haha. I should be doing hw but....I suck at that. I never got much of a sufficient answer for the last question I posted, so Im wonderng if I should just close it and ask a different one.

OpenStudy (anonymous):

thats what works for me... keep nagging until you get results

OpenStudy (anonymous):

Ill probably just close my current question and ask a new one. Otherwise Im just on here helpng out until im ready to sleep xD

OpenStudy (anonymous):

brilliance should be shared :)

OpenStudy (anonymous):

I hate to think of myself that way. Im just done a lot of math and have been able to understand a lot of it. I still have to study just like anyone else and work hard. Trig Ive just had enough practice with over time.

OpenStudy (anonymous):

just take a compliment!

OpenStudy (anonymous):

Fine xD Thanks :P

OpenStudy (anonymous):

ok so I need help again.. I dont know if my brain is processing less since I am getting more and more tired. 2cos^2(theta)-2cos(theta)-1=0 I thought it would be 120 and 240

OpenStudy (anonymous):

Oh, now we switched to the same question but cosine. So it's going to be the same thing as before. We had to do quadratic formula to get results. And since theyre the SAME coefficients, the quadratic formula will give the same results. So youcan actually take the same answers we got from the previous question, the whole 1/2 +/- sqrt{3}/2 and take inverse cosine of those instead of inverse sine like we did last time.

OpenStudy (anonymous):

h=-16t^2+vtsin(theta) Give the equation for the height, if v is 1,500 feet per second and θ is 30°. h = -16t2+ t

OpenStudy (anonymous):

so how exactly would I insert it... would I just insert -16t^2+1500tsin(pi/3)

OpenStudy (anonymous):

UNfortunately got all those stupid glitchy question marks, so not sure what Im missing. But yeah, I'd say plug in just as you did. And Ill assume it wants you to solve for t afterward but yeah, that looks fine :)

OpenStudy (anonymous):

ok I am stuck again... ok so here goes one problem... cos(A-85 deg)=sqrt(3)/2

OpenStudy (anonymous):

so I knew it would be 30 and 330.... so I added 85 6o 30.. and got 115 + 360. then I added 85 to 330 so that it would be 55 + 360...

OpenStudy (anonymous):

Right, so 115 + 360k and 55 + 360k, that works :3

OpenStudy (anonymous):

oh i know what i did wrong...

OpenStudy (anonymous):

thanks....

OpenStudy (anonymous):

Was there something wrong with that answer?

OpenStudy (anonymous):

i didnt insert it right... so I am stuck on two more problems then I get to go to sleep

OpenStudy (anonymous):

Lol, cool cool.

OpenStudy (anonymous):

i am officially on the last problem of this assignment

OpenStudy (anonymous):

w00t, lol. Hopefully not a hard one.

OpenStudy (anonymous):

im done!!! yay!! i only have 4 more assignments!!! :(

OpenStudy (anonymous):

For the rest of thesemester or what?

OpenStudy (anonymous):

noooo I have to turn in tomorrow

OpenStudy (anonymous):

Eeek, that far behind? o.o

OpenStudy (anonymous):

yes!! and Im sick... please tell me you will be here tomorrow... you are the smarted guy here

OpenStudy (anonymous):

Ill be here. I cant say when Ill be on, but I will be.